# PriRoAgg: Achieving Robust Model Aggregation with Minimum Privacy Leakage for Federated Learning

### v1

@@ Abstract
Federated learning (FL) has recently gained significant momentum due to its potential to leverage large-scale distributed user data while preserving user privacy. However, the typical paradigm of FL faces challenges of both privacy and robustness: the transmitted model updates can potentially leak sensitive user information, and the lack of central control of the local training process leaves the global model susceptible to malicious manipulations on model updates. Current solutions attempting to address both problems under the one-server FL setting fall short in the following aspects: 1) designed for simple validity checks that are insufficient against advanced attacks (e.g., checking norm of individual update); and 2) partial privacy leakage for more complicated robust aggregation algorithms (e.g., distances between model updates are leaked for multi-Krum). In this work, we formalize a novel security notion of aggregated privacy that characterizes the minimum amount of user information, in the form of some aggregated statistics of users’ updates, that is necessary to be revealed to accomplish more advanced robust aggregation. We develop a general framework PriRoAgg, utilizing Lagrange coded computing and distributed zero-knowledge proof, to execute a wide range of robust aggregation algorithms while satisfying aggregated privacy. As concrete instantiations of PriRoAgg, we construct two secure and robust protocols based on state-of-the-art robust algorithms, for which we provide full theoretical analyses on security and complexity. Extensive experiments are conducted for these protocols, demonstrating their robustness against various model integrity attacks, and their efficiency advantages over baselines.
@@ I Introduction
At the core of PriRoAgg are three key components: (1) Robust algorithms; (2) Secret sharing with Lagrange coded computing (LCC) [37]; and (3) Zero-knowledge proof (zk-proof). This framework is designed to incorporate robust algorithms that can compute a reliable final aggregation against potential malicious updates. Users’ private data remains in perfect secrecy within a certain number of colluding users since it is secret shared via LCC. In the presence of Byzantine users who may arbitrarily deviate from the protocol, PriRoAgg further leverages SNIP [38], a secret sharing-based zk-proof, to ensure the consistency of local executions. Users generate proofs that are verifiable by the collaboration of all other users and the server. The server can orchestrate the users to perform the final model aggregation securely and robustly.
@@
  • We significantly improve the communication efficiency of PriRoAgg, utilizing the constant-size commitment scheme [33, 39] and extending the zk-proof scheme to LCC.
@@ II-B Privacy-preserving Model Aggregation
Contradicting the early concept, studies have revealed that conventional federated learning poses risks of leaking users’ privacy. Works such as [2, 4, 42, 5] have shown that by launching model inversion attacks, it is possible to reveal information about the local training data of individual users from their local updates or gradients. To address this issue, secure aggregation is introduced to resolve the aggregation problem without leaking any information except for the final aggregation result [43, 30, 19, 20, 18, 21, 22, 23, 24].
The main approaches in secure aggregation include Secret sharing (SS) [32, 33, 30, 31], Differential privacy (DP) [43, 18, 21, 22] and Homomorphic encryption (HE) [25, 26, 27]. Secret sharing enables a user to split its local update (secret) into multiple secret shares, where the secret remains theoretically private as long as no entity holds more shares than a specified threshold. Shareholders (other users) can collectively aggregate these secret shares to compute the aggregated secret. The server collects the aggregation of shares and reconstructs the actual aggregation of local updates. The differential privacy technique is to add controlled noise or randomness to updates before aggregation. This ensures that the aggregated results do not disclose sensitive information about local training data. Homomorphic encryption is a powerful cryptographic technique that allows computations to be performed on encrypted data without decrypting it. The server can exploit such a feature to perform a secure aggregation and decrypt the real aggregation without privacy leakage.
@@ II-C Robust Model Aggregation
Aside from concerns for privacy, the issue of model robustness has been pervasively discussed in FL. The defense strategies and attack methods have been competitively growing in recent years. Generally, there are two types of attacks of interest. Some malicious users can launch untargeted attack, during which they upload arbitrary false models to impede the server from completing the model aggregation correctly [44, 45, 5]. There are also omniscient untargeted attacks where the adversary is aware of other benign updates, so it can design the update to launch stronger attacks [46]. The other type is targeted attack, where the malicious users deliberately implant the global model with a backdoor without affecting the main learning task  [47, 48, 49, 50, 51, 52, 53].
To detect such fraudulent model updates, robust FL aggregation has been introduced by many studies. At the onset of the introduction of model poisoning, a few noise-filtering-based algorithms are put forward to select honest ones across users [9, 10, 11, 12, 13]. The statistical traits of local updates are exploited to capture the malformed ones and compute a final aggregation without them. However, malicious users can manipulate the updates to bypass the defense so that get selected for the final aggregation [54, 46]. Competitively, Some recent robust aggregation protocols against poisoning attacks have been proposed to address this issue [14, 15, 16, 17].
In addition, various backdoor defense methods have been proposed to counter deliberate manipulations on local updates, as evidenced by studies such as [55, 56, 57, 52, 58, 59, 60, 61, 62, 63, 59, 64]. In [62], the coordinate-wise median and trimmed mean techniques are employed to achieve promising performance in robust aggregation. Several studies also investigate the impact of the sign of updated gradients [63, 59, 64]. Particularly, [63] is the first to utilize majority voting on gradient signs during final aggregation to enhance robustness and convergence. Subsequent works investigate how to use the coordinate-wise sign information to guide the final aggregation; for instance, [59] uses the sign to adjust the learning rate in each round to defend against backdoor attacks, while SignGuard  [64] clusters the signs to filter malicious gradients.
@@ II-D Robust and Privacy-preserving Model Aggregation
Recent works attempt to simultaneously address both aspects of privacy and robustness [32, 43, 21, 33, 30, 65, 35, 31]. BREA [32] is the first FL protocol known to consider privacy-preserving and robustness simultaneously. It employs the statistical algorithm multi-Krum [14] to protect the final aggregation from Byzantine local updates, and uses secret sharing to hide the updates from the adversary. A state-of-the-art protocol, EIFFeL  [30], provides an exemplary solution that addresses robustness and privacy simultaneously. It significantly improves the privacy guarantees to the privacy using new cryptography preliminaries. Regarding robustness, EIFFeL shows that users can use any function to validate the integrity of their individual updates. EIFFeL does not require users to upload the local updates for integrity validation, instead, a zk-proof for correctly performing the evaluation on updates. RoFL [31] further extends the norm bounding validation in EIFFeL and demonstrates different norm constraints also mitigate the effect of backdoor attacks. It also boosts the efficiency of the protocol in addition. A recent work, ByITFL, proposes a theoretically secure protocol [28] that provides robustness in update aggregation. It builds upon the Byzantine-resilient protocol FLTrust [66] to further provide privacy, especially in scenarios involving colluding users alongside a curious server. Other protocols have also been proposed to offer privacy and robustness by exploiting multiple-server architectures, such as [34, 35, 36].
@@ II-E Challenges and Motivation
Current robust and secure protocols face limitations in either privacy-preserving or robust aggregation performance. The primary issue of BREA [32] is the incomplete privacy left with security breaches. The multi-Krum algorithm [14] used in the protocol reveals the pairwise distances between the local updates to the server. Such pairwise distances potentially provide information to a curious server to reconstruct users’ local updates. On the other hand, for works that provide provable privacy, the defensive measurements are insufficient to advance adversary attacks. EIFFeL [30] and RoFL [31] perform the validation function on user updates individually; hence, their defense cannot leverage the overall statistical information across users. For example, malicious users can easily bypass detection by keeping their updates below the norm threshold while still implanting a backdoor.
For general robust aggregations, cryptography techniques mentioned in Section II-B have limitations in different aspects. Though HE provably addresses privacy concerns, a critical limitation is its impractical complexity [25, 26, 27, 34, 28]. In practice, the computational cost is $10^{3}$ to $10^{4}$ times higher than the non-encrypted vanilla FL, even when the models used are as simple as logistic regression and multi-layer perceptron. Hence, many HE protocols suffer from high complexity and become highly impractical when sophisticated models are involved. DP alleviates the high complexity associated with HE, providing better efficiency along with provably secure solution [43, 18, 21, 22]. However, a major issue with DP is that its random masking hinders the detection of malicious updates. A malicious update potentially becomes indistinguishable to many robust algorithms by masquerading as a benign one using a masked update. How to evaluate and provide robustness against deliberately manipulated updates still remains an open problem in DP [29, 22].
@@ III Problem Formulation
We note that the semi-honest (honest-but-curious) server setting is pervasively considered in previous works on secure and robust federated learning [32, 43, 23, 69, 34, 70, 36]. In comparison with works [32, 33, 71, 69, 34, 36] which consider the coexistence of semi-honest server(s) and malicious users but acting separately, we also consider the possibility of server-user collusion. For example, an adversary can manipulate the messages sent from a portion of users, such that the aggregated model update, observed at the curious server, reveals more information about the other victim users’ private data. Also, for the single-server setting considered in this work, dealing with a Byzantine server who can arbitrarily deviate from the protocol is extremely challenging. Even for the protocols in [30] and [31] that claim to be secure against a malicious server actively inferring users’ private data, they are vulnerable to recently developed active attacks for data leakage (see, e.g., [72]). Hence, these protocols are essentially secure against a semi-honest server.
For a set of user updates $\{\mathbf{x}_{i}|i\in[N]\}$, we denote a robust algorithm by $\Omega$ and its output by $\mathbf{x}_{\Omega}$, i.e., $\mathbf{x}_{\Omega}\leftarrow\Omega(\{\mathbf{x}_{i}|i\in[N]\})$. For example, [14] computes $\mathbf{x}_{\Omega}$ by minimizing the distances between $\mathbf{x}_{\Omega}$ and $\mathbf{x}_{i}$s; and [73] computes $\mathbf{x}_{\Omega}$ by taking the geometric median of $\mathbf{x}_{i}$s. Given an aggregation algorithm, the goal is to compute $\mathbf{x}_{\Omega}$ with the minimum information leakage to the adversary. For secure aggregation, where $\Omega$ represents summation, protocols are designed to compute the sum of $\mathbf{x}_{i}$s as the final result such that nothing else is exposed to the adversary. However, for more complicated $\Omega$ (e.g., [14, 74, 75]), additional information is often leaked to the curious server and users to accomplish the aggregation. To formalize the minimal possible amount of information leaked to the adversary, we generalize the privacy guarantee to a notion called aggregated privacy, which requires that no additional information is exposed except the summation of some function of $\mathbf{x}_{i}$. Specifically, to compute $\mathbf{x}_{\Omega}$ in a privacy-preserving manner, we slightly relax the security requirement compared with secure aggregation, by allowing the adversary to know nothing but some aggregated statistics of local updates. We make this notion more concrete in the following definition:
@@
Given users’ local updates $\{\mathbf{x}_{i},\forall i\in[N]\}$, a robust aggregation algorithm $\Omega$, and some function $\phi_{\Omega}(\cdot)$ on an individual update, an aggregation protocol satisfies aggregated privacy if nothing other than the aggregation $\sum_{i\in[N]}\phi_{\Omega}(\mathbf{x}_{i})$ is revealed to the adversary.
Remark: As a special instance, when $\Omega$ stands for summation, $\phi_{\Omega}$ is the identity function, i.e., $\sum_{i\in[N]}\phi_{\Omega}(\mathbf{x}_{i})=\sum_{i\in[N]}\mathbf{x}_{i}$, then the aggregated privacy reduces to the same privacy constraint as in concventional secure aggregation. Achieving aggregated privacy for general $\Omega$ is highly non-trivial. For example, BREA [32] and ByzSecAgg [33] utilize the Krum algorithm, where the pairwise distances between $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are revealed to the curious server, hence violating the aggregated privacy. Also, in EIFFeL [30] and RoFL [31], a valid predicate is applied on updates, where the validity of $\mathbf{x}_{i}$ is revealed to all parties, which also violates the aggregated privacy.
To design provably secure robust protocols with aggregated privacy, we formulate a secure multi-party computation (MPC) problem with $N$ input updates $\{\mathbf{x}_{i}|i\in[N]\}$ from $N$ FL users, denoted as ${SRA}_{\Omega}$ (short for secure and robust aggregation with $\Omega$). The adversary gains full control over a subset $U_{M}$ of corrupted users and can access the local state of the server; it runs any next-message algorithm denoted by $\mathcal{A}$.
@@
Given a robust aggregation algorithm $\Omega$ and inputs $\{\mathbf{x}_{i}|i\in[N]\}$, a protocol $\Pi_{\Omega}$ securely computes $\Omega$ with aggregated privacy if it satisfies:
  • Integrity. Consider the output of the protocol execution $\mathbf{x}_{agg}=\Pi_{\Omega}(\{\mathbf{x}_{i}|i\in[N]\})$, and $\mathbf{x}_{\Omega}=\Omega(\{\mathbf{x}_{i}|i\in[N]\})$. Then, there always exists a proper security parameter $\kappa$, such that the following holds:
@@
  • Privacy. For any adversary running $\mathcal{A}$ who controls a subset $U_{M}$ of malicious users and a semi-honest server, and $U_{H}=[N]\backslash U_{M}$ denotes the rest of honest users, there exists a probabilistic polynomial-time (P.P.T.) simulator $\mathcal{S}$, such that
@@
    Secret-shared non-interactive proof [38] provides a distributed zk-proof for the authenticity of the users’ operations. We use this technique to guarantee that the users have honestly performed the function evaluations on their shares.

### v2

@@ Abstract
Federated learning (FL), as a promising machine learning paradigm for large-scale distributed data, faces two security challenges of privacy and robustness: the transmitted model updates potentially leak sensitive user information, and the lack of central control over local model updates leaves the global model susceptible to malicious attacks. Current solutions attempting to address both problems under the one-server FL setting fall short in the following aspects: 1) design for simple validity checks that are insufficient against advanced attacks (e.g., checking norm of individual update); and 2) have partial privacy leakage for more complicated robust aggregation algorithms (e.g., distances between model updates are leaked for multi-Krum). In this work, we formalize a novel security notion of aggregated privacy that characterizes the minimum amount of user information, in the form of aggregated statistics of users’ updates, that is necessary to be revealed to accomplish more advanced robust aggregation. We develop a general framework PriRoAgg, utilizing Lagrange coded computing and distributed zero-knowledge proof, to execute a wide range of robust aggregation algorithms while satisfying aggregated privacy. As concrete instantiations of PriRoAgg, we construct two secure and robust protocols based on state-of-the-art robust algorithms, for which we provide full theoretical analyses on security and complexity. Extensive experiments are conducted for these protocols, demonstrating their robustness against various model integrity attacks, and their efficiency advantages over baselines.
@@ I Introduction
At the core of PriRoAgg are three key components: (1) Robust algorithms; (2) Secret sharing with Lagrange coded computing (LCC) [37]; and (3) Zero-knowledge proof (ZKP). This framework is designed to incorporate robust algorithms that can compute a reliable final aggregation against potential malicious updates. Users’ private data remains in perfect secrecy within a certain number of colluding users via LCC. In the presence of Byzantine users who may arbitrarily deviate from the protocol, PriRoAgg further leverages SNIP [38], a secret sharing-based ZKP, to ensure the consistency of local executions. Users generate proofs that are verifiable by the collaboration of all other users and the server. The server can orchestrate the users to perform the final model aggregation securely and robustly.
@@
  • We significantly improve the communication efficiency of PriRoAgg, utilizing the constant-size commitment scheme [33, 39] and extending the ZKP scheme to LCC.
@@ II-B Privacy-preserving Model Aggregation
Contradicting the early concept of FL’s privacy, studies have revealed that conventional federated learning poses risks of leaking users’ data. Works such as [2, 4, 42, 5, 43] have shown that by launching gradient inversion attacks, it is possible to reveal information about the local training data from individual users’ local updates. To address this issue, secure aggregation is introduced to protect individual information during aggregation process and only reveal the final aggregated model  [44, 19, 20, 18, 21, 22, 23, 24]. Aiming at preserving privacy, the commonly assumed threat model is honest-but-curious (semi-honest), where all parties follow the protocol but try as much as possible to infer user’s individual information. Moreover, to mitigate the risk of the server manipulating the aggregation results, several subsequent secure aggregation approaches have evolved into verifiable secure aggregation methods to provide model integrity [45, 46].
Secret sharing (SS) is a typical way to achieve secure aggregation [32, 33, 30, 31, 47]. It enables a user to split its local update (secret) into multiple secret shares, where the secret remains theoretically private as long as no entity holds more shares than a specified threshold. Shareholders (other users) can collectively aggregate these secret shares to compute the aggregated secret. The server collects the aggregation of shares and reconstructs the actual aggregation of local updates. Differential privacy (DP) is also popularly used in secure aggregation protocols  [44, 18, 21, 22, 48]. This technique is to add controlled noise or randomness to updates before aggregation. This ensures that the aggregated results do not disclose sensitive information about local training data. Homomorphic encryption (HE) is another powerful cryptographic technique that allows computations to be performed on encrypted data without decrypting it, therefore, exploited in secure aggregation protocols  [25, 26, 27, 45]. The server can exploit such a feature to perform a secure aggregation and decrypt the real aggregation without privacy leakage. Zero-knowledge proof as a common MPC technique, is used by a few verifiable aggregation protocols to provide verifiability to the model aggregations performed on the server side[49, 45, 46, 47].
@@ II-C Robust Model Aggregation
Aside from concerns for privacy, the issue of model robustness has been pervasively discussed in FL. It refers to attacks launched by malicious users who aim to sabotage the learning process from convergence or implement a backdoor to the global model, which impair the model robustness overall. Generally, there are two types of attacks of interest. Malicious users can launch untargeted attack, during which they upload arbitrary false models to impede the server from completing the model aggregation correctly [50, 51, 5]. There are also omniscient untargeted attacks where the adversary is aware of other benign updates, so it can design the update to launch stronger attacks [52] or evade the gradient sifting algorithms [53]. The other type is targeted attack or backdoor attack, where the malicious users try to inject a backdoor to the global model without affecting the main learning task  [54, 55, 56, 57, 58, 59, 60].
To detect such fraudulent user updates, robust FL aggregation has been introduced by an abundance of studies. At the onset of the introduction of model poisoning, a few noise-filtering-based algorithms are put forward to select honest ones across users [9, 10, 11, 12, 13]. The statistical traits of local updates are exploited to capture the malformed ones and compute a final aggregation without them. However, new attacks manipulate the updates to bypass the defense so that get selected for the final aggregation [61, 52]. Competitively, Some recent robust aggregation protocols against poisoning attacks have been proposed to address this issue [14, 15, 16, 17].
Backdoor attacks are more stealthy and harder to detect as it requires similar update statistics to preserve the main task performance. Various backdoor defense methods have been proposed to counter deliberate manipulations on local updates, as evidenced by studies such as [62, 63, 64, 59, 65, 66, 67, 68, 69, 70, 66, 71]. In [69], the coordinate-wise median and trimmed mean techniques are employed to achieve promising performance in robust aggregation. Several studies notice the impact of the sign differences of honest and malicious gradients [70, 66, 71]. Particularly, [70] is the first to utilize majority voting on gradient signs during final aggregation to enhance robustness and convergence. Subsequent works investigate how to use the coordinate-wise sign information to guide the final aggregation; for instance, [66] uses the sign to adjust the learning rate in each round to defend against backdoor attacks, while SignGuard  [71] clusters the signs to filter malicious gradients.
@@ II-D Robust and Privacy-preserving Model Aggregation
Simultaneously addressing both aspects in II-B and II-C is crucial in real-world applications; however, it presents significant challenges. Several studies have endeavored to tackle this issue [32, 21, 33, 30, 35, 31, 34, 36, 48, 47, 72]. BREA [32] is the first privacy-preserving and robust FL protocol known, who uses secret sharing scheme to hide local updates, and multi-Krum algorithm to remove malicious updates  [14]. [33] significantly improves on the computational and communication efficiency of BREA. [21] investigates how local and central DP alleviates various attacks and offers a quantifiable framework for the trade-off between privacy and robustness. A privacy-preserving protocol, EIFFeL  [30], provides a secret-sharing based secure aggregation solution while provide robustness by well-formedness verification on user updates. It does not require users to upload the local updates for integrity validation, instead, using a ZKP scheme for correctly performing norm integrity check locally. RoFL [31] further extends the norm bounding validation in EIFFeL and demonstrates different norm constraints also mitigate the effect of backdoor attacks. It boosts the efficiency of the protocol in addition. ByITFL [28] proposes a theoretically secure protocol that provides robustness by update norm validation as well. It is built upon the Byzantine-resilient protocol FLTrust [73] to further provide privacy by LCC, especially in scenarios involving colluding users alongside a curious server. Trust-ZKFL [47] explores a protocol based on ZKFL [49] that provides user privacy, aggregation robustness and model integrity under a non-colluding threat model. It proposes a rank-based scheme and a similarity-based scheme to compute inter-user information to robustly guide the final model aggregation. [72] proposes a privacy-preserving and robust protocol by introducing a group of trusted autonomous auditing entities to perform filtering for malicious encrypted updates. It utilizes additive HE and a consensus scheme to relay the encrypted messages processed by the auditors. Other protocols offer privacy and robustness guarantees by exploiting multiple-server architectures, such as [34, 35, 36, 48].
@@ II-E Challenges and Motivation
Current robust and secure protocols face limitations in either (a) privacy-preserving guarantee or (b) robust aggregation performance. (a) On one hand, the primary issue in  [32, 33, 74] is the privacy leakage of individual user information through intermediate computation results in plaintext. In BREA[32] and ByzSecAgg[33], the multi-Krum algorithm [14] is used, which reveals the pairwise distances between the local updates. Such pairwise distances potentially provide information to a curious server to reconstruct users’ local updates. In Trust-ZKFL [47], the coordinate-wise distances between user updates are also leaked to the attacker under collusion. In rank-based scheme or similarity-based scheme, for two users in comparison, when one of the users colludes with the server or one user from the reconstruction subset, the corrupted party is able to retrieve another user’s local model by simple computations. (b) On the other hand, works that provide provable privacy under user-server collusion are unable to defend against malicious user attacks. Specifically, EIFFeL [30], RoFL [31] and [72] perform validation functions on user updates individually; hence, their defense cannot leverage the overall statistical information across users. For example, malicious users can easily bypass detection by keeping their updates below the norm threshold while still implanting a backdoor.
For general robust aggregations, cryptography techniques mentioned in Section II-B have limitations in different aspects. Though HE provably addresses privacy concerns, a critical limitation is its impractical complexity [25, 26, 27, 34, 28, 72]. In practice, the computational cost is $10^{3}$ to $10^{4}$ times greater than that of the non-encrypted vanilla FL, and the models under consideration are restricted to simple logistic regression models in [34, 72]. Hence, HE protocols suffer from extremely high complexity, rendering them highly impractical when sophisticated models are involved. Compared with HE, DP alleviates the high complexity and provides comparable efficiency with vanilla FL, while give provably secure solutions [44, 18, 21, 22]. However, the major issue with DP is the random masking that hinders the detection of malicious updates. A malicious update potentially becomes indistinguishable to many robust algorithms by masquerading as a benign one using a masked update. How to evaluate and provide robustness against deliberately manipulated updates in DP still remains an open problem [29, 22].
@@ III Problem Formulation
We note that the honest-but-curious (semi-honest) server setting is pervasively considered in previous works on secure and robust federated learning [32, 33, 44, 23, 77, 34, 48, 36, 47, 72]. In comparison with works like  [78, 77, 34, 36, 47, 48, 72] who consider the coexistence of semi-honest server(s) and malicious users but acting separately, we also consider the possibility of server-user collusion. That is, an adversary can manipulate the messages sent from the corrupted users, such that the aggregated model update, observed at the curious server, reveals more information about the other victim users’ private data. Also, for the single-server setting considered in this work, dealing with a Byzantine server who can arbitrarily deviate from the protocol is extremely challenging. Even for the protocols in [30] and [31] that claim to be secure against a malicious server actively inferring users’ private data, they are vulnerable to recently developed active attacks for data leakage (see, e.g., [79]). Hence, these protocols are essentially secure against a semi-honest server.
For a set of user updates $\{\mathbf{x}_{i}|i\in[N]\}$, we denote a robust algorithm by $\Omega$ and its output by $\mathbf{x}_{\Omega}$, i.e., $\mathbf{x}_{\Omega}\leftarrow\Omega(\{\mathbf{x}_{i}|i\in[N]\})$. For example, [14] optimizes $\mathbf{x}_{\Omega}$ by minimizing the distances between $\mathbf{x}_{\Omega}$ and $\mathbf{x}_{i}$s; and [80] computes $\mathbf{x}_{\Omega}$ by finding the geometric median of $\mathbf{x}_{i}$s. Generally for any robust model aggregation, it can be represented by defining a $\Omega$. Then given an aggregation algorithm, the goal is to compute $\mathbf{x}_{\Omega}$ with the minimum information leakage to the adversary. For secure aggregation, where $\Omega$ represents update summation, protocols are designed to compute $\Sigma\mathbf{x}_{i}$ as the final result such that nothing else is exposed. However, when $\Omega$ is complicated for a robust aggregation (e.g., [14, 81, 74, 66]), additional information is often leaked to the curious server and users. To formalize the minimal possible amount of information leaked to the adversary, we generalize the privacy guarantee to a notion called aggregated privacy, which requires that no additional information is exposed, except the summation of a given function of $\mathbf{x}_{i}$. Therefore, to compute $\mathbf{x}_{\Omega}$ with a sophisticated $\Omega$ in a privacy-preserving manner, we slightly relax the security requirement in secure aggregation to certain aggregated statistics of local updates, but not allowing any information leakage related to any specific user. We formalize this notion in the following definition:
@@
Given users’ local updates $\{\mathbf{x}_{i},\forall i\in[N]\}$, a robust aggregation algorithm $\Omega$, and the function $\phi_{\Omega}(\cdot)$ on an individual update, an aggregation protocol satisfies aggregated privacy if nothing other than the aggregation $\sum_{i\in[N]}\phi_{\Omega}(\mathbf{x}_{i})$ is revealed to the adversary.
Remark 1: As a special instance, when $\Omega$ stands for a plain summation and $\phi_{\Omega}$ is the identity function, i.e., $\sum_{i\in[N]}\phi_{\Omega}(\mathbf{x}_{i})=\sum_{i\in[N]}\mathbf{x}_{i}$, the aggregated privacy reduces to the conventional privacy constraint in secure aggregation. However, plain summation of individual updates does not provide any robustness to malicious updates.
Remark 2: Achieving aggregated privacy for robust aggregation algorithm $\Omega$ is highly non-trivial, and none of previous works have fulfilled such definition. BREA [32] and ByzSecAgg [33] utilize the Krum algorithm, where the pairwise distances between $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are revealed to the curious server, hence violating the aggregated privacy. Similarly, Trust-ZKFL [47] reveals coordinate-wise distance or cosine similarity between user updates, to not only the server but also a subset of users. In EIFFeL [30] and RoFL [31], a validation predicate, typically a norm bound, is applied to individual updates, revealing the validity of $\mathbf{x}_{i}$ to all parties. This approach violates aggregated privacy. To mitigate the privacy leakage of individual users, [72] introduces a trusted and autonomous auditing system but still reveals the validity of $\mathbf{x}_{i}$ to all.
To design provably secure robust protocols with aggregated privacy, we formulate a secure multi-party computation (MPC) problem with $N$ input updates $\{\mathbf{x}_{i}|i\in[N]\}$ from $N$ FL users, denoting the MPC problem as ${SRA}_{\Omega}$ (short for secure and robust aggregation with $\Omega$). The adversary controls over a corrupted subset $U_{M}$ of users and can access the local state of the server; it runs any next-message algorithm, $\mathcal{A}$.
@@
Given a robust aggregation algorithm $\Omega$ and inputs $\{\mathbf{x}_{i}|i\in[N]\}$, a protocol $\Pi_{\Omega}$ securely executes $\Omega$ satisfying aggregated privacy if it satisfies:
  • Integrity. Consider the output of the protocol execution $\mathbf{x}_{agg}=\Pi_{\Omega}(\{\mathbf{x}_{i}|i\in[N]\})$, and the plaintext execution $\mathbf{x}_{\Omega}=\Omega(\{\mathbf{x}_{i}|i\in[N]\})$. There always exists a proper security parameter $\kappa$, such that the following holds:
@@
  • Privacy. For any adversary running $\mathcal{A}$ who controls a malicious subset $U_{M}$ of users and a semi-honest server, and $U_{H}=[N]\backslash U_{M}$ denotes the rest of honest users, there exists a probabilistic polynomial-time (P.P.T.) simulator $\mathcal{S}$, such that
@@
    Secret-shared non-interactive proof [38] provides a distributed ZKP for the authenticity of the users’ operations. We use this technique to guarantee that the users have honestly performed the function evaluations on their shares.